# Li Zhou1, Lutong Yu1, Dongchu Xie1, Shaohuan Cheng3, Wenyan Li4, Haizhou Li1, 2

### v2

@@ Appendix B Prompts for VQA
"answer": The selected correct option (e.g., "A", "B", "C", or "D").
@@
"answer": The selected correct option (e.g., "A", "B", "C", or "D").
@@ C.1 Caption Extraction and Caption Edit Prompts
Transform this traditional Hanfu description: "{caption}" into a modern clothing design by strategically incorporating its key elements. Your task is to: 1. Choose ONE specific modern garment type as the base (select from: hoodie, blazer, casual wear, sportswear, trench coat, streetwear, or business attire) 2. Identify 3-5 distinctive elements from the original Hanfu description (such as collar style, sleeve design, waist details, fabric patterns, or color schemes) 3. Describe exactly how these Hanfu elements are integrated into the modern garment: - WHERE each element is placed on the modern garment - HOW each element is adapted to suit contemporary fashion - WHY these particular elements were chosen (cultural significance) 4. Ensure the final design: - Is primarily a modern, wearable garment for everyday contexts - Clearly displays its Hanfu inspiration through intentional design choices - Balances contemporary style with traditional Chinese aesthetics - Appeals to modern fashion sensibilities while honoring cultural heritage The output should provide a detailed description of this hybrid garment, suitable for image generation, focusing on both visual appearance and construction details.
@@ C.4 Evaluation Feedback on Generated Images
Evaluator 3: Most generated garments were identifiable as functional clothing. However, the criterion of "compatibility with contemporary lifestyles" was deemed ambiguous, as interpretations and acceptance thresholds vary substantially across individuals.

### v3

@@ Appendix B Prompts for VQA
”answer”: The selected correct option (e.g., ”A”, ”B”, ”C”, or ”D”).
@@
”answer”: The selected correct option (e.g., ”A”, ”B”, ”C”, or ”D”).
@@ C.1 Caption Extraction and Caption Edit Prompts
Transform this traditional Hanfu description: ”{caption}” into a modern clothing design by strategically incorporating its key elements. Your task is to: 1. Choose ONE specific modern garment type as the base (select from: hoodie, blazer, casual wear, sportswear, trench coat, streetwear, or business attire) 2. Identify 3-5 distinctive elements from the original Hanfu description (such as collar style, sleeve design, waist details, fabric patterns, or color schemes) 3. Describe exactly how these Hanfu elements are integrated into the modern garment: - WHERE each element is placed on the modern garment - HOW each element is adapted to suit contemporary fashion - WHY these particular elements were chosen (cultural significance) 4. Ensure the final design: - Is primarily a modern, wearable garment for everyday contexts - Clearly displays its Hanfu inspiration through intentional design choices - Balances contemporary style with traditional Chinese aesthetics - Appeals to modern fashion sensibilities while honoring cultural heritage The output should provide a detailed description of this hybrid garment, suitable for image generation, focusing on both visual appearance and construction details.
@@ C.4 Evaluation Feedback on Generated Images
Evaluator 3: Most generated garments were identifiable as functional clothing. However, the criterion of ”compatibility with contemporary lifestyles” was deemed ambiguous, as interpretations and acceptance thresholds vary substantially across individuals.